Find the LCM and HCF of the following integers by applying the prime factorisation method.
 8, 9 and 25

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

8, 9 and 25
Writing the product of prime factors for all the three numbers, we get,
8=2×2×2×1
9=3×3×1
25=5×5×1
Therefore,
HCF(8,9,25)=1
LCM(8,9,25) =2×2×2×3×3×5×5= 1800

​Thanks